This file documents all modifications and changes to DEIMOS
which are noteworthy.
CCD 1A and 1B video signal coax swap [2003-Aug-19]
Having determined that the SMA/SMB swap appeared to fix
the intermittent dropped bit problem, Bob had Hector swap the
video signal coaxes, thus introducing a second swap which has
the effect of undoing the SMA/SMB swap of two weeks ago. The
next effect is that the output images are back to normal. See
this document for further details.
CCD1 SMA/SMB connector swap [2003-Aug-05]
We have been having occasional problems with dropped bits on CCD
1A. To diagnose the source, Bob Kibrick swapped the SMA/SMB
connectors on CCD1. This will swap the data in the resulting
Ne lamp replaced [2003-Jul-30]
Replaced dim neon lamp.
DEIMOS Guider Shutter Repair [2003-May-07]
Replaced the guider shutter with p/n 2000734. Of the three
shutters that were tested this is the only one that worked like
shutters should. The voltage that opens the shutter is 14vdc,
however, the voltage fluctuates anywhere from 8vdc to 14vdc.
More tests are needed on the shutters that have the star shape
Dewar Filter WHeel Repair [2003-Jan-07]
Filter wheel quit rotating during a reconfig. Removed panels and
found the motor spur gear slipped on the motor shaft to
disengage the filter wheel. Repositioned the spur gear, R&R the
set screw to apply Lock-Tite blue, and reinit the stage. Stage
appears to be working ok.
Cleaned Slitmask Cassette [2002-Oct-31]
Rodriguez and Wagner performed the following service as requested
except for using air. Dust was excessive and it was felt that
contaminating the optics was highly probable.
Recommend modification to stage to allow easy removal of
cassette for cleaning outside of instrument. Modification
would involve changing blind bolts to accessable bolts where
cassette attaches to actuator plate.
Replaced Failed Guider Shutter [2003-Oct-07]
Removed the filter wheel to inspect the shutter. Verified that
the shutter was opening partially and needed to be replaced or
secured open. After inspecting the filter wheel assembly, the
decision was made to remove the shutter. In order for this to
happen, one of the two filter wheel plates needed to be removed.
Remove the filter wheel plate closest to the filter wheel.
Removed and replaced the old shutter (p/n 04l214-6v; s/n
9913243) with a new shutter (p/n 04l214-6v; s/n 9913245). This
was the ESI spare shutter. Needed to replace the 6vdc cable in
order to install the DEIMOS connector. Verified that the new
shutter operated properly before installing on to the camera
head. The filter wheel plate that was removed was not put back
into the instrument. This should not effect any filter wheel or
guider camera operations, and will facilitate replacing shutters
in the future. Mounted the filter wheel and verified proper
rotation. Installed camera access cover.
Install Pull-Up Resistors to Cure LED Problem [2002-Sep-24/25]
Partial Implemented ECR 255. Verified TV filter stage and TV
focus had the pull-up resistors installed. Modified Grating
select stage , slitmask stage , and Slitmask cassette
environment sense control box. with 1K pull-up
resistors. Verified operation. On-going.
Completed implementation of ECR 255. Installed pull-up
resistors in the CCD focus and translatiion stages,
filterwheel stage, grating tilt #3 stage, and grating tilt #4
stage. Functional check good.
Slitmask Cassette Repair [2002-Sep-13]
Troubleshoot DEIMOS problem with slitmask juke box jamming. We
weren't able to move slit mask assembly. Found grating cover
was missing a screw and dismounted from its normal position.
Screw was found inside instrument. The cover was then caught
against juke box during moves. The grating cover was bent from
the incident. The cover was bent back and reinstalled properly
with locktite. Slitmask mechanism was then moved mechanically
and verified to be operational. Instrument Specialist and
Astronomers are informed of the situation. Pictures of grating
cover forwarded to G. Wirth.
During troubleshooting, metal filings were removed from the
brass guide rollers in front of the slit box mechanism. We
also found leaky glycol fitting at the guide camera.
Installed teflon on fitting and verified no leaks.
Guider Repair [2002-Aug-11]
Continued from previous day
The key was that we discovered that one of the boards in the
head, which we found out today was the serial clock board, could
be swapped with the spare even though they were different
versions. It made sense to swap this since the original error
was a serial clock error. We didn't do it Saturday because we
didn't know if we could swap different version boards. Once the
swap was made, the camera worked.
In order to install the camera we discovered the filter wheel
is outfitted with a quick disconnect so the filters can be
kept safe while the filter wheel mechanism goes in. After the
camera was all hooked back up together, turned on the glycol
flow and took good images. Sandy and Mason aligned the
camera. It is in proper focus when viewing from the front the
camera is rotated to the maximum clockwise position that the
slop in tne mounting bolts allow.
Other Stuff to be aware of that we did and checked --
The RBC PXL head is being brought down to be sent back for repair. Miguel has
- removed air so that we could manually move hatch up and down. air
- removed hatch b. replaced hatch b.
- removed front two lower quarter panels on snout with guider camera oriented
horizontal on lower half of snout. panels were replaced.
- covered window with low lint optical cloth. removed cloth.
- the hatch covers around the PXL CEU2 controller were removed. hatch covers
- DEIMOS was undefined, then later defined.
Guider Repair [2002-Aug-10]
DEIMOS Guider is NOT functional. The guider will not be
available for the rest of the DEIMOS run. Efforts should be
made to form a backup observing plan.
Today's efforts were focused on removing the DEIMOS guider
camera head and trying to get it or the spare head to
work. We were unsuccessful.
The basic status is that last night's troubleshooting was
aimed at isolating the failure. It was isolated to the
DEIMOS PXL camera head. Because the camera head alignment
was very difficult, the goal was to try to repair the
DEIMOS camera head and put it back in with as little
disrupt to alignment as possible.
A VERY arduous process was undertaken to get the camera
head removed so we could inspect and swap boards. Thanks
to Dave Cowley for the procedure. I did get a few pictures
that should help. The following steps were taken --
This took about 4+ hours.
- drove tel up and over to NRT2/RT4 position
- backed DEIMOS up to get access to front
- rotated so hatch was horizontal on lower half of disk
- locked DEIMOS
- removed lower 2 arc panels on snout
- removed unused side B hatch (dropped a washer in snout, didn't realize it
wasn't tapped, could not locate washer, probably fell thru grating)
- Tim put low lint cloth across back of window and cardboard to protect it.
- Disconnected compressed air and bled lines in prep for removing 4" beam in
front of snout.
- Taped hatch closed.
- Turned out we didn't need to remove beam to get camera and filter wheel
out. Tim did need to fabricate two cut-off allen wrenches, though.
- Unplugged filter wheel cables
- Unplugged camera head cables except DB25 power cable which was screwed on.
- Released mounting bolts from front (have pictures) and camera head and
filter wheel with their respective mounting plates came out together. This
step very awkward and difficult and somewhat risky. The filter wheel was
separated from the camera head and stored in safe location. The camera head
was still attached by the power cable and glycol lines, Removed power cable.
- shutoff glycol to PXL
- removed PXL QDs on cooling lines
- unrouted PXL head cables to give enough slack to test camera outside side
Once the camera head was out we proceeded to
troubleshoot. There are 4 boards in the camera head. 3 of
the boards are easily accessible, so we tried swapping
them, one at a time, with the RBC guider camera head. This
did not work -- same serial clock error. In order to try
the fourth board, we had to de-mount the camera head from
it's mounting plate.
When we got the fourth board out we realized they were
different versions. Decided not to take a chance and
jeopardize the CCD. Instead decided with Sandy Faber to
try and put the RBC camera back together (what we were
calling the spare) and try it in DEIMOS.
Configured RBC back to original configuration. Connected
to DEIMOS controller thru DEIMOS cables and got a green
light on the controller. Tried to take and image but got
an error in kcamtest that pointed to a controller prom
needing to be swapped. Swapped the PROM. This time no
At this point encountered a major irritation. Could not
take images with kcamtest on huakini and couldn't get
exceed to work from a PC. It works all the way up until
you read out an image then it crashes. WE NEED TO FIX
Ran xguide instead since we knew PROM was
compatible. Image read out but image had all zeroes. This
same configuration was tried last night during
troubleshooting and worked so something has failed between
then and today. Most obvious thing was the video cable
from head to controller since we know we read all zeroes
when it becomes unplugged. Inspected and buzzed cable. All
ok. Tried a spare set of cables, still all zeroes.
It could be that during the swapping of camera head boards
that one of the boards were damaged or the boards got
mixed up. Opened up the spare camera head and double
checked serial numbers again...all correct as we had
assumed. Last gasp was to try the video board from the
DEIMOS head in the spare head. This time we did not get
all zeroes, but we did get an awful looking image with big
stripes of all zeros and stripes of a few hundred
counts. Anyway totally not functional. Saved a cam image
for later analysis.
At this point Sandy/Greg decided to observe without the guider.
We've put the front and rear hatches and panels back. Both
guider heads and filter wheel are removed and in the
instrument lab. The inside of the snout was checked and
double checked. The low lint cloth was removed from the
window. Air is back on, hatch works, PXL glycol left
off. DEIMOS is defined, fill complete. Glenn is in the
process of changing a grating for tonites modified
We did not have time to put metal tape along all the seams
for light tightness along the hatch covers.
The following things were taken to the instrument lab --
- Both guider camera heads. (I will ask Glenn to cart
them down tomorrow assuming we have a less busy day tomorrow.)
- The filterwheel assembly
- guider mounting plates for RBC and DEIMOS
- cables and hardware
Huge thanks to Bill Healy, he's been here with me since
9am. Thanks also to Saloga for helping demount the camera
and to Glenn for helping out and staying late. Thanks to
Jason, Tom, Steve and Glenn for staying late last night to
troubleshoot. Sorry to the observers, I really don't think
we could have done much more.